As a 2.0-kg object moves from (11i-5j)m to (7i-8j)m, the constant resultant force acting on it is equal to (4i-3j)N. If the speed of the object at the initial position is 4.0 m/s, what is its kinetic energy at its final position?
Answer #1

m Mass of object = 2.0 kg Force acting on object = F = (42 - 3 f) N 11111 (432443,2 = 116+9 = 125 = 5N The deject moved fromKinetic Energy at its final position a / mn2 (K.E.) final a 4 x 2x 41 (K.E.)c = 41 J